OpenStudy (jack1):

please help... finding this Q kinda difficult:

OpenStudy (jack1):

in need to turn \[\large f(s) = \frac 1{s(s+6)(s+10)(s+1)}\] into partial fractions

OpenStudy (jack1):

I'm pretty sure I go from here to : \[\large f(s) = \frac A{s} + \frac B{(s+6)} + \frac C{(s+10)} + \frac D{(s+1)} \]... yeah?

OpenStudy (jack1):

then cross multiply and expand out the brackets to get a common denominator...? maybe?

OpenStudy (raden):

yeah, i think so

OpenStudy (jack1):

cheers @RadEn

OpenStudy (raden):

i got A = 1/60 for the others, i got 3 equations : B+C+D = -1/60 11B + 7C + 16D = -17/60 10B + 6C + 60D = -76/60 obvious, we can solve them (b,c,and d) by elimination and subt methods. good luck :)